# Greedy Algorithm for Structured Bandits: A Sharp Characterization of Asymptotic Success / Failure*

### v2

@@ Appendix A $\mathtt{StructuredCB}$ with tie-breaking
$E_{1}$ ={|$=\left\{\,\absolutevalue{\bar{r}_{{\mathtt{warm}}}(x,a)-f_{\mathtt{dec}}(x,a)}<\Gamma(f_{\mathtt{dec}})/2\quad\text{for each $x\in\mathcal{X}$ and arm $a\notin\mathcal{A}^{*}_{\mathtt{dec}}(x)$}\,\right\},$
$E_{2}$ ={|$=\left\{\,\absolutevalue{\bar{r}_{t}(x,\pi_{\mathtt{dec}}(x))-f^{*}(x,\pi_{\mathtt{dec}}(x))}<\Gamma(f_{\mathtt{dec}})/2\quad\text{for each $x\in\mathcal{X},a\in\mathcal{A}^{*}_{\mathtt{dec}}(x)$, and round $t>T_{0}$}\,\right\}.$
@@ Appendix B Novelty of self-identifiability
We argue that self-identifiability is a novel notion. Specifically, we compare it to (i) knowing the optimal value, and (ii) Graves-Lai coefficient being 00.
First, one could ask if self-identifiability is equivalent to knowing the value of the best arm. However, the former does not imply the latter. Consider the simple example =$\mathcal{F}=\left\{(3,1),(2,1)\right\}$. Both functions are self-identifiable in $\mathcal{F}$, but clearly the optimal value differs.
@@
Now, one could ask if self-identifiability is equivalent to $\mathsf{GLC}(\mathcal{M},M)=0$. We observe that this is not the case: the two notions are incomparable. For a counterexample, consider $\mathtt{StructuredMAB}$ with two arms and =$\mathcal{F}=\left\{(2,1),(0.5,1)\right\}$. Problem instance $f^{*}=(0.5,1)$ is self-identifiable, since revealing the sub-optimal arm as having reward 0.5 immediately rules out $(2,1)$ as being the true model. But the $\mathsf{GLC}>0$, since to ascertain $(0.5,1)$ as being the true model one necessarily has to choose the 1st arm and experiment. On the other hand, one can see $f^{*}=(2,1)$ is not self-identifiable but has $\mathsf{GLC}=0$. In this example, $\mathtt{Greedy}$ succeeds when $\mathsf{GLC}>0$ (larger $\mathsf{GLC}$ suggests larger regret of the optimal algorithm in $\mathsf{GLC}$-based theory) but fails when $\mathsf{GLC}=0$ (lower $\mathsf{GLC}$ suggests lower regret of the optimal algorithm in $\mathsf{GLC}$-based theory)! Hence $\mathsf{GLC}$ does not capture the per-instance behavior of $\mathtt{Greedy}$.
@@
We complete the proof of Theorem 3.3(a) as follows. The regret incurred during the warmup data is at most $T_{0}$. Fix any round $t>T_{0}$. After the warmup data, we know with probability $1-1/t$, any suboptimal arm can be pulled at most $(K/\Gamma(f^{*})^{2})\cdot O(\log t)$ times after the warmup data, and the regret is $(K/\Gamma(f^{*})^{2})\cdot O(\log t)$. With the remaining probability $1/t$, the regret is at most $O(t)$. Hence, the theorem follows.
@@
$E_{1}=\left\{\,\absolutevalue{\bar{r}_{{\mathtt{warm}}}(a)-f_{\mathtt{dec}}(a)}<\Gamma(f_{\mathtt{dec}})/2\quad\text{for each arm $a\neq a_{\mathtt{dec}}$}\,\right\}.$
$E_{2}=\left\{\,\forall t>T_{0},\;\absolutevalue{\bar{r}_{t}(a_{\mathtt{dec}})-f^{*}(a_{\mathtt{dec}})}\leq\Gamma(f_{\mathtt{dec}})/2\,\right\}.$
@@
For any $x$ in the interval $[f_{\mathtt{dec}}(a)-\Gamma(f_{\mathtt{dec}})/2,f_{\mathtt{dec}}(a)+\Gamma(f_{\mathtt{dec}})/2]$, by boundedness of mean reward, we have
|$\absolutevalue{x-f^{*}(a)}\leq 2.$
Then, the density of $x$ at any point on the interval $[f_{\mathtt{dec}}(a)-\Gamma(f_{\mathtt{dec}})/2,f_{\mathtt{dec}}(a)+\Gamma(f_{\mathtt{dec}})/2]$ is at least
@@
$\Pr[\absolutevalue{\bar{r}_{{\mathtt{warm}}}(a)-f_{\mathtt{dec}}(a)}\leq\Gamma(f_{\mathtt{dec}})/2]$ $\geq\frac{\Gamma(f_{\mathtt{dec}})}{\sqrt{2\pi\sigma^{2}}}\exp(-2/\sigma^{2}).$
@@
$E_{3}=\left\{\exists t>T_{0},\absolutevalue{\bar{r}_{t}-f_{\mathtt{dec}}(a_{\mathtt{dec}})}>\Gamma(f_{\mathtt{dec}})/2\right\},$
@@
$E_{1}$ ={|$=\left\{\,\absolutevalue{\bar{r}_{{\mathtt{warm}}}(x,a)-f^{\dagger}(x,a)}<\Gamma(f^{\dagger})/2\quad\text{for each $x\in\mathcal{X}$ and arm $a\neq\pi^{\dagger}(x)$}\,\right\},$ (D.1)
$E_{2}$ ={|$=\left\{\,\absolutevalue{\bar{r}_{t}(x,\pi^{\dagger}(x))-f^{*}(x,\pi^{\dagger}(x))}<\Gamma(f^{\dagger})/2\quad\text{for each $x\in\mathcal{X}$ and round $t>T_{0}$}\,\right\}.$ (D.2)
@@
$\mathtt{MSE}_{s}(f)\geq T^{\prime}\cdot(\Gamma(f^{*})/2)^{2}\geq\absolutevalue{\mathcal{X}}K\cdot O(\log(\absolutevalue{\mathcal{X}}Kt)).$
@@
Conditional on the event in Lemma D.1, the expected total number of times of suboptimal policy execution is no larger than |$\absolutevalue{\mathcal{X}}KT^{\prime}/p_{0}=(\absolutevalue{\mathcal{X}}K/\Gamma(f^{*}))^{2}/p_{0}\cdot O(\log t)$.
@@
Since we have that deterministically $M_{t}<\absolutevalue{\mathcal{X}}KT^{\prime}$, we know that the total number of times of suboptimal policy execution $N_{t}=\sum_{s=1}^{t}I_{s}$ satisfies
$\operatornamewithlimits{\mathbb{E}}[N_{t}]=\operatornamewithlimits{\mathbb{E}}[\sum_{s=1}^{t}I_{s}]\leq\operatornamewithlimits{\mathbb{E}}[\sum_{s=1}^{t}(\operatornamewithlimits{\mathbb{E}}[M_{s}\mid M_{s-1}]-M_{s-1})]/p_{0}<\absolutevalue{\mathcal{X}}KT^{\prime}/p_{0}.$
@@
Proof of Theorem 4.3(a). The regret incurred in the warmup phase is at most $T_{0}$. With probability $1-1/t$, the number of suboptimal policy pulls can be bounded as in the lemma above. With the remaining $1/t$ probability the regret is at most $O(t)$. Finally, by the regret decomposition lemma (Lemma 4.5 in Lattimore and Szepesvári (2020)), we have
$\operatornamewithlimits{\mathbb{E}}[R(t)]$ $\leq T_{0}+\absolutevalue{\mathcal{X}}KT^{\prime}/p_{0}+O(1)$
$\leq T_{0}+(\absolutevalue{\mathcal{X}}K/\Gamma(f^{*}))^{2}/p_{0}\cdot O(\log t)$
@@
$E_{1}$ ={|$=\left\{\,\absolutevalue{\bar{r}_{{\mathtt{warm}}}(x,a)-f_{\mathtt{dec}}(x,a)}<\Gamma(f_{\mathtt{dec}})/2\quad\text{for each $x\in\mathcal{X}$ and arm $a\neq\pi_{\mathtt{dec}}(x)$}\,\right\},$
$E_{2}$ ={|$=\left\{\,\absolutevalue{\bar{r}_{t}(x,\pi_{\mathtt{dec}}(x))-f^{*}(x,\pi_{\mathtt{dec}}(x))}<\Gamma(f_{\mathtt{dec}})/2\quad\text{for each $x\in\mathcal{X}$ and round $t>T_{0}$}\,\right\}.$
@@
Event $E_{1}$ happens with probability at least $\Omega(\frac{\Gamma(f_{\mathtt{dec}})\exp(-2/\sigma^{2})}{\sigma})^{\absolutevalue{\mathcal{X}}K}$.
@@
$E_{3}=\left\{\exists t,x,\absolutevalue{\bar{r}_{t}(x,\pi_{\mathtt{dec}}(x))-f_{\mathtt{dec}}(x,\pi_{\mathtt{dec}}(x))}\geq\Gamma(f_{\mathtt{dec}})/2\right\}$
@@
$\Pr[E_{1}\cap E_{2}]\geq\left[\log(|\mathcal{X}|)\exp(-2(\log\absolutevalue{\mathcal{X}})^{2}/\Gamma(f_{\mathtt{dec}}))^{2}\right]^{\absolutevalue{\mathcal{X}}K}.$
@@
[…]
@@
Choose $\delta=1/(t|\Pi||\mathcal{M}|)$ and that $\log(1/\delta)=O(|\Pi|t)$, then with probability $1-1/t$, the total number of suboptimal policies executions can be upper bounded by
@@
|$\absolutevalue{\bar{r}_{\tau}(a)-f^{*}(a)}<\varepsilon/2.$
Take any reward vector $f^{\prime}$ such that |$\absolutevalue{f(a^{\prime})-f(a)}>\varepsilon$. Then we have
|$\absolutevalue{\bar{r}_{\tau}(a)-f(a)}\geq\varepsilon/2.$
@@
Therefore any $f^{\prime}$ with |$\absolutevalue{f^{\prime}(a)-f(a)}\geq\varepsilon$ cannot possibly be minimizing $\mathtt{MSE}_{\tau}(\cdot)$. That is to say, the reward vector $f_{\tau}$ minimizing $\mathtt{MSE}_{\tau}(\cdot)$ must have |$\absolutevalue{f_{\tau}(a)-f^{*}(a)}<\varepsilon$. Then, by the strong notion of $\varepsilon$-self-identifiability, we precisely know that arm $a$ is also a suboptimal arm for the reward vector $f_{\tau}$. Hence, we obtain a contradiction, and arm $a$ cannot possibly be chosen this round.
@@
$E_{1}=\left\{\,\absolutevalue{\bar{r}_{{\mathtt{warm}}}(a)-(f_{\mathtt{dec}}(a)-\varepsilon/(2\sqrt{K}))}<\varepsilon/(4\sqrt{K})\quad\text{for each arm $a\neq a_{\mathtt{dec}}$}\,\right\},$
$E_{2}=\left\{\,\forall t>T_{0},\;\absolutevalue{\bar{r}_{t}(a_{\mathtt{dec}})-f^{*}(a_{\mathtt{dec}})}\leq\varepsilon/(4\sqrt{K})\,\right\}.$
These events mirror (3.2) and (3.3), but with two changes: (1) we replace the confidence radius $\Gamma(f_{\mathtt{dec}})/2$ by $\varepsilon/(4\sqrt{K})$, and (2) shift the baseline value of $f_{\mathtt{dec}}(a)$ by $-\varepsilon/(2\sqrt{K})$ in $E_{1}$.
@@
For any $x$ in the interval $[f_{\mathtt{dec}}(a)-3\varepsilon/(4\sqrt{K}),f_{\mathtt{dec}}(a)-\varepsilon/(4\sqrt{K})]$, by boundedness of mean reward, we have
|$\absolutevalue{x-f^{*}(a)}\leq 2.$
Then, the density of $x$ at any point on the interval $[f_{\mathtt{dec}}(a)-3\varepsilon/(4\sqrt{K}),f_{\mathtt{dec}}(a)-\varepsilon/(4\sqrt{K})]$ is at least
@@
$\Pr[\absolutevalue{\bar{r}_{{\mathtt{warm}}}(a)-(f_{\mathtt{dec}}(a)-\varepsilon/(2\sqrt{K})}<\varepsilon/(4\sqrt{K})]$ $\geq\frac{\varepsilon}{2\sqrt{2\pi\sigma^{2}K}}\exp(-2/\sigma^{2}).$
@@
$E_{3}=\left\{\exists t>T_{0},\absolutevalue{\bar{r}_{t}-f_{\mathtt{dec}}(a_{\mathtt{dec}})}>\varepsilon/(4\sqrt{K})\right\},$

### v3

@@ Appendix A $\mathtt{StructuredCB}$ with tie-breaking
[…]
@@ Appendix B Novelty of self-identifiability
We argue that self-identifiability is a novel notion. Specifically, we compare it to (i) knowing the optimal value, and (ii) Graves-Lai coefficient being 0.
First, one could ask if self-identifiability is equivalent to knowing the value of the best arm. However, the former does not imply the latter. Consider the simple example $\mathcal{F}=\left\{(3,1),(2,1)\right\}$. Both functions are self-identifiable in $\mathcal{F}$, but clearly the optimal value differs.
@@
Now, one could ask if self-identifiability is equivalent to $\mathsf{GLC}(\mathcal{M},M)=0$. We observe that this is not the case: the two notions are incomparable. For a counterexample, consider $\mathtt{StructuredMAB}$ with two arms and $\mathcal{F}=\left\{(2,1),(0.5,1)\right\}$. Problem instance $f^{*}=(0.5,1)$ is self-identifiable, since revealing the sub-optimal arm as having reward 0.5 immediately rules out $(2,1)$ as being the true model. But the $\mathsf{GLC}>0$, since to ascertain $(0.5,1)$ as being the true model one necessarily has to choose the 1st arm and experiment. On the other hand, one can see $f^{*}=(2,1)$ is not self-identifiable but has $\mathsf{GLC}=0$. In this example, $\mathtt{Greedy}$ succeeds when $\mathsf{GLC}>0$ (larger $\mathsf{GLC}$ suggests larger regret of the optimal algorithm in $\mathsf{GLC}$-based theory) but fails when $\mathsf{GLC}=0$ (lower $\mathsf{GLC}$ suggests lower regret of the optimal algorithm in $\mathsf{GLC}$-based theory)! Hence $\mathsf{GLC}$ does not capture the per-instance behavior of $\mathtt{Greedy}$.
@@
We complete the proof of Theorem 3.3(a) as follows. The regret incurred during the warmup data is at most $T_{0}$. Fix any round $t>T_{0}$. After the warmup data, we know with probability 11/t1-1/t, any suboptimal arm can be pulled at most $(K/\Gamma(f^{*})^{2})\cdot O(\log t)$ times after the warmup data, and the regret is $(K/\Gamma(f^{*})^{2})\cdot O(\log t)$. With the remaining probability $1/t$, the regret is at most $O(t)$. Hence, the theorem follows.
@@
[…]
@@
For any $x$ in the interval $f_{\mathtt{dec}}(a)-\Gamma(f_{\mathtt{dec}})/2,f_{\mathtt{dec}}(a)+\Gamma(f_{\mathtt{dec}})/2]$, by boundedness of mean reward, we have
[…]
Then, the density of $x$ at any point on the interval $f_{\mathtt{dec}}(a)-\Gamma(f_{\mathtt{dec}})/2,f_{\mathtt{dec}}(a)+\Gamma(f_{\mathtt{dec}})/2]$ is at least
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Proof of Theorem 4.3(a). The regret incurred in the warmup phase is at most $T_{0}$. With probability 11/t1-1/t, the number of suboptimal policy pulls can be bounded as in the lemma above. With the remaining $1/t$ probability the regret is at most $O(t)$. Finally, by the regret decomposition lemma (Lemma 4.5 in Lattimore and Szepesvári (2020)), we have
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Let $\pi^{\circ}$ be any suboptimal policy. Fix $\delta\in\left(\,0,\frac{1}{|\mathcal{M}|}\,\right)$. With probability at least $1-|\mathcal{M}|\delta$, the policy $\pi^{\circ}$ can be executed for at most $O\left(\,\sigma_{0}^{2}\,\Gamma^{-2}\,\ln(1/\delta)\,\right)$ rounds.
@@
Choose $\delta=1/(t|\Pi||\mathcal{M}|)$ and that $\log(1/\delta)=O(|\Pi|t)$, then with probability 11/t1-1/t, the total number of suboptimal policies executions can be upper bounded by
@@
[…]
@@
[…]
@@
[…]
These events mirror (3.2) and (3.3), but with two changes: (1) we replace the confidence radius $\Gamma(f_{\mathtt{dec}})/2$ by $\varepsilon/(4\sqrt{K})$, and (2) shift the baseline value of $f_{\mathtt{dec}}(a)$ by $\varepsilon/(2\sqrt{K})$ in $E_{1}$.
@@
For any $x$ in the interval $f_{\mathtt{dec}}(a)-3\varepsilon/(4\sqrt{K}),f_{\mathtt{dec}}(a)-\varepsilon/(4\sqrt{K})]$, by boundedness of mean reward, we have
[…]
Then, the density of $x$ at any point on the interval $f_{\mathtt{dec}}(a)-3\varepsilon/(4\sqrt{K}),f_{\mathtt{dec}}(a)-\varepsilon/(4\sqrt{K})]$ is at least
@@
[…]
@@
[…]